OpenStudy (anonymous):

if (3,1) and (-3,5) are my points how do i convert them into a standard form equasion?

OpenStudy (anonymous):

for this problem yo would have to find the slope of this line. Do you know how to do that?

OpenStudy (anonymous):

yes the slope is -4/3

OpenStudy (anonymous):

Nah, it's -4/6

OpenStudy (anonymous):

im sorry -2/3 i dont know what i was thinking

OpenStudy (anonymous):

it happens :)

OpenStudy (anonymous):

i simplified

OpenStudy (anonymous):

now what

OpenStudy (anonymous):

now can you, using the slope you have, write in point slope form?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

y=mx+b y=-2/3 x+b

OpenStudy (anonymous):

do you know y-y1=m(x-x1)

OpenStudy (anonymous):

choose one of the points, plug in y and x accordingly, and solve for b

OpenStudy (anonymous):

ah then i put the (3,1) or (-3,5) in for x and y then i have my y intercept then im dont. ok thankyou i get it now

OpenStudy (anonymous):

ur welcome! good job!

OpenStudy (anonymous):

no you are not done, you wanted standard form not point slope form

OpenStudy (anonymous):

do you know how to change from point slope form to standard form

OpenStudy (anonymous):

I thought y=mx+b was standard form ;(

OpenStudy (anonymous):

thats what im thinking

OpenStudy (anonymous):

no, standard form is AX+BY=C

OpenStudy (anonymous):

wo

OpenStudy (anonymous):

Yeah, not everything in math is cut and dry

OpenStudy (anonymous):

are you sure

OpenStudy (anonymous):

what are the a,b,and c?

OpenStudy (anonymous):

yea i am sure, you said you wanted standard form right?

OpenStudy (anonymous):

yes i know how to swich them i just thought they were each other

OpenStudy (anonymous):

okay then